OpenStudy (sgayton27):

What is the radius of a copper sphere (density = 8.92 g/cm^3) whose mass is 3.75 g? The volume of a sphere is given by the equation V=(4/3)(pi)r^3. Please help me get this one started!

OpenStudy (aaronq):

the formula for density is: \(\sf density=\dfrac{mass}{volume}\)

OpenStudy (aaronq):

So plug stuff in: \(\sf 8.92 ~g/cm^3=\dfrac{3.75~ g}{\dfrac{4}{3}\pi r^3}\) solve for r.